OpenStudy (anonymous):

Express x^2+6x-1 in the form f(x)=a(x-h)^2+k. Express x^2+6x-1 in the form f(x)=a(x-h)^2+k. @Mathematics

OpenStudy (mathmagician):

\[x^2+6x-1=x^2+6x+9-9-1=(x+3)^2-10\]

OpenStudy (mathmagician):

or (x-(-3))^2-10

OpenStudy (anonymous):

Where did you get the 9 and -9 from?

OpenStudy (mathmagician):

9-9 is equal to 0, so you can add and subtract 9, and i did it because 9 is required for a full sqare formula- \[x^2+6x+9=(x+3)^2\]

OpenStudy (amistre64):

that is the whole purpose of this exercise, to get you comfortable with completing the square

OpenStudy (amistre64):

the other means is to just go with the axis of symmetry and try to weave it back in

OpenStudy (anonymous):

Completing the Square is: \[\frac{-b}{2a} \] Correct so then I would do what: (-3)^2+6(-3)-1=

OpenStudy (amistre64):

f(x)=a(x-h)^2+k f(x) = a(x^2-2hx+h^2) + k f(x) = ax^2-2hax+ah^2 + k g(x) = mx^2+nx+p then equate f and g, woulda used a b and c but they seemed to be used lol

OpenStudy (anonymous):

or do I add the 9 to both sides?

OpenStudy (amistre64):

f(x) = x^2+6x-1 ; looks that a = 1 soo f(x) = x^2-2hx+h^2 + k 2hx = 6x 2h = 6 h = 6/2 = 3

OpenStudy (amistre64):

f(x) = x^2-6x+3^2 + k 3^2 + k = -1 9 + k = -1 k = -10

OpenStudy (amistre64):

f(x)=a(x-h)^2+k f(x)=(x-3)^2 -10

OpenStudy (anonymous):

ahh... \[(\frac{-b}{2a})^{2} = (\frac{-6}{2})^2= (-3)^2 = 9\]

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

instead of adding 9 to both sides; just add 0 to one side

OpenStudy (amistre64):

9-9 = 0

OpenStudy (anonymous):

ahh ok got it.

OpenStudy (anonymous):

Thanks you for the assistance.

OpenStudy (anonymous):

it's -10 cause it's the opposite? or it should equal -10 for a reason?

OpenStudy (anonymous):

ahh my bad got it now.